# Arbitrary Entropy Policy Optimization: Entropy Is Controllable in Reinforcement Fine-tuning

### v4

@@ Abstract
Reinforcement fine-tuning (RFT) is essential for enhancing the reasoning capabilities of large language models (LLM), yet the widely adopted Group Relative Policy Optimization (GRPO) suffers from entropy collapse, where entropy monotonically decreases, exploration vanishes, and policies converge prematurely. Existing entropy-regularized methods only partially alleviate this issue while introducing bias and instability, leaving entropy control unresolved and the connection between entropy, exploration, and performance unclear. We propose Arbitrary Entropy Policy Optimization (AEPO), which eliminates entropy collapse by replacing entropy bonuses with REINFORCE policy gradient on temperature-adjusted distributions and stabilizing entropy through temperature regulation. AEPO integrates three key designs: policy gradient as regularization, distribution as regularization, and REINFORCE as regularization, enabling precise entropy control without distorting optimization. Experiments demonstrate three major contributions: AEPO (1) stabilizes entropy at arbitrary target levels, effectively removing collapse in GRPO; (2) reveals a non-monotonic relation where performance first improves then declines with increasing entropy, clarifying the link between entropy, exploration, and reasoning; and (3) generalizes beyond entropy, providing a broader RFT paradigm where superior target distributions can serve as REINFORCE regularizers. Our code is available at https://github.com/597358816/AEPO.
@@
Figure 1: Entropy across five runs of AEPO. By adjusting only the parameter $\mathcal{H}$, entropy can be stably maintained at different levels.
@@ 1 Introduction
Reinforcement fine-tuning (RFT) has become a cornerstone for enhancing the reasoning capabilities of large language models (LLM) (GLM et al., 2024; Touvron et al., 2023; Schulman et al., 2017; Rafailov et al., 2023; Zhong et al., 2024; Wang et al., 2024). Among existing approaches, Group Relative Policy Optimization (GRPO) has gained wide adoption due to its efficiency and scalability (Shao et al., 2024; Liu et al., 2024; Guo et al., 2025). However, GRPO suffers from a well-documented drawback: entropy collapse: As training progresses, policy entropy declines monotonically, sampled outputs converge to nearly identical solutions, and the model prematurely adopts a deterministic policy with limited exploration (Yu et al., 2025; Li et al., 2025a; Zhang et al., 2025). This severely restricts the ability of RFT to discover diverse reasoning strategies. Existing remedies typically rely on entropy regularization and clip refine, which inevitably introduce bias and instability into the optimization objective, creating a trade-off between exploration and stability (Hou et al., 2025; Cui et al., 2025; Cheng et al., 2025; Shen, 2025). EFRame is the first to achieve large-scale entropy control, but it remains a qualitative analysis without conducting a quantitative study of entropy (Wang et al., 2025a).
@@
Although the issue of entropy collapse in GRPO has been repeatedly noted, it remains unresolved: there is still no principled algorithm capable of precisely regulating entropy throughout training. Existing approaches often treat entropy only as a side indicator—for instance, using it to split steps or guide auxiliary heuristics—while the policy entropy itself continues to collapse as training progresses (Wang et al., 2025b; Zheng et al., 2025; Li et al., 2025b; Cheng et al., 2025). As a result, the relationship among entropy, exploration, and performance remains unclear: it is uncertain whether entropy is a sufficient proxy for exploration or whether exploration itself consistently improves training outcomes. If significant performance improvement can be observed by adjusting entropy to a better range, it would indicate that exploration plays a crucial role in this process. If arbitrary entropy control can be achieved during RFT, it would make it possible to realize exploration at any desired degree and, in turn, establish a principled connection among entropy, exploration, and performance.
Motivated by these challenges, we propose Arbitrary Entropy Policy Optimization (AEPO), a novel policy gradient that directly addresses entropy collapse. AEPO replaces conventional entropy bonuses with a REINFORCE policy gradient (Williams, 1992; Sutton et al., 1999) applied to samples drawn from temperature-adjusted distributions. Entropy is further stabilized through temperature regulation, ensuring that batch-level entropy remains oscillating around an arbitrary constant $\mathcal{H}$ throughout training, as shown in Fig. 1. AEPO achieves entropy control through three key design components:
  • Policy gradient as regularization: Instead of relying on a conventional entropy bonus, we employ a full policy gradient term applied to samples that naturally exhibit high- or low-entropy properties. This design ensures that entropy never dominates the optimization objective, allowing the model to monotonically explore toward higher accuracy throughout training.
  • Distribution as regularization: Entropy is regulated via temperature-adjusted distributions. Given a predefined entropy threshold $\mathcal{H}$, when the previous step’s entropy $\mathcal{H}(\pi_{\theta_{\text{old}}})<\mathcal{H}$, we regard the distribution under $T_{\text{high}}>1$ as a better candidate and mix in a small proportion of its samples. Conversely, when $\mathcal{H}(\pi_{\theta_{\text{old}}})\geq\mathcal{H}$, we instead sample from the distribution under $T_{\text{low}}<1$.
  • REINFORCE as regularization: In Reinforcement Learning with Verifiable Reward (RLVR), the REINFORCE algorithm can filter out negative samples in an unbiased manner, allowing positive samples to form a unidirectional gradient toward a better distribution. This guides the model to optimize from positive samples that align with the target distribution. If negative samples were included at this stage, the entropy control mechanism would fail.
@@
  • Controllable entropy: We propose Arbitrary Entropy Policy Optimization that can stabilize entropy at arbitrary target levels, effectively eliminating entropy collapse in GRPO.
@@
  • Generalizability: Beyond entropy control, AEPO provides a broader paradigm for RFT. When a target distribution $\pi^{*}$ is identified as superior to the current policy $\pi_{\theta}$, samples from $\pi^{*}$ can be used to construct a REINFORCE policy gradient as a regularization term, which allows $\pi$ to progressively approximate $\pi^{*}$ during long-horizon training.
@@ 2 Related Work
RFT has become a central paradigm for post-training large language models (LLMs), aligning them with human feedback and task-specific objectives. Early methods such as RLHF leveraged policy optimization (e.g., PPO) to encode human preferences (OpenAI, 2023; Team et al., 2024; Wei et al., 2023; Liu et al., 2023), while Direct Preference Optimization (DPO) (Rafailov et al., 2024) later improved efficiency by optimizing policies directly from preference data. Recent models like DeepSeek-R1 (Guo et al., 2025) and Kimi-1.5 (Team et al., 2025) extend RFT through hybrid reward formulations and scalable optimization. Among them, Group Relative Policy Optimization (GRPO) (Shao et al., 2024; Liu et al., 2024) has become the de facto baseline for reasoning-focused RFT, yet suffers from entropy collapse that limits exploration of diverse reasoning strategies.
Entropy has long been regarded as a proxy for exploration in reinforcement optimization. Classical methods employ entropy regularization to stabilize training and encourage diversity (Sutton et al., 1999; Williams, 1992). More recent studies extend this idea to GRPO by introducing entropy bonuses into rewards or advantages (Cheng et al., 2025; Cui et al., 2025; Shen, 2025). However, such approaches only yield coarse-grained effects: entropy still collapses as training proceeds, or the added bias destabilizes optimization. EFRame (Wang et al., 2025a) represents the first framework to achieve large-scale entropy control through conditional sampling and replay, yet its treatment remains qualitative and lacks quantitative characterization of the entropy–exploration–performance relationship.
@@ 3 Preliminary
Our work focuses on fine-tuning LLM using Reinforcement Learning (RL) for tasks with verifiable solutions, such as mathematical reasoning and code generation.
@@ 3.2 Entropy-regularization variants
Entropy-Adv(Cheng et al., 2025) proposed an entropy-guided advantage shaping method. The key idea is to inject an entropy-based term into the advantage function during policy optimization. They define an entropy-based advantage term $\psi(\mathcal{H}_{t})$ and use it to shape advantage:
@@
$\nabla_{\theta}\mathcal{J}^{\text{shaped}}(\theta)=$ $\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(O|q)},$
@@
Table 1: Comparison of entropy-based optimization objectives.
Policy Gradient
Entropy-Reg $\sum_{t}(\hat{A}_{t}\cdot\nabla_{\theta}\text{log}\pi_{\theta}(o_{t}|q,o_{<t})+\lambda\cdot\nabla_{\theta}\mathcal{H}_{t})$
Entropy-Adv $\sum_{t}A_{i,t}^{\text{shaped}}\cdot\nabla_{\theta}\text{log}\pi_{\theta}(o_{t}|q,o_{<t})$
@@ 4 Method
Arbitrary Entropy Policy Optimization (AEPO) is designed to achieve precise and stable control of policy entropy during RFT. Unlike conventional entropy-regularized methods, AEPO does not introduce explicit entropy bonuses; instead, it adjusts the training distribution and policy gradient to regulate entropy implicitly yet controllably. This section first introduces the key premises underlying AEPO’s design, which reveal the relationship between temperature, entropy, and policy updates, and then details the algorithmic formulation of AEPO.
$\mathcal{J}_{GRPO}(\theta)=\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,\hat{A}_{i,t},\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,\hat{A}_{i,t}\Big],$ (1)
$\mathcal{J}_{AEPO}(\theta)=\;\mathcal{J}_{GRPO}(\theta)+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi^{T}_{\theta_{\text{old}}}(O|q)}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,R(q,o_{i}),\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,R(q,o_{i})\Big],$
@@
4.1 Premise
The design of AEPO is built upon two empirical premises that connect temperature-based sampling to entropy dynamics. These premises establish the foundation for controllable entropy modulation without introducing bias into the optimization objective.
Premise 1.
@@ 4.1 Premise
Previous studies, such as Du et al. (2025), show that increasing the sampling temperature generally broadens the model’s output distribution and raises its entropy. Although the correspondence between higher temperature and higher entropy is not a universally proven principle, it holds—or approximately holds—under most intuitive and commonly used model settings. In the context of RFT, temperature thus serves as a practical external variable for indirectly regulating entropy, enabling controllable entropy adjustment.
Figure 2: Entropy dynamics under temperature-controlled sampling. High-temperature positive samples increase entropy, promoting exploration, while low-temperature positive samples reduce entropy, leading to more deterministic behavior.
Lemma 4.1.
(Cui et al., 2025) If the actor policy $\pi_{\theta}$ is a tabular softmax policy updated via natural policy gradient (Kakade, 2001) with step size $\eta$. Then the change in policy entropy between two steps approximately satisfies:
$\begin{split}\mathcal{H}(\pi_{\theta}^{k+1})-\mathcal{H}(\pi_{\theta}^{k})\approx&-\eta\cdot\mathbb{E}_{s\sim d_{\mu}^{k}}\text{Cov}_{a\sim\pi^{k}_{\theta}(\cdot\mid s)}\\ &\!\bigl[\log\!\pi^{k}_{\theta}(a\mid s),\,A^{\pi^{k}}(s,a)\bigr].\end{split}$
The proof can be seen in Liu (2025) and Cui et al. (2025). $\mathcal{H}$ indicates the policy entropy of policy model, and Cov denotes covariance, $\pi^{k}_{\theta}$ is the policy at step $k$, and $A^{\pi^{k}}(s,a)$ is the advantage function of action $a$ under state $s$. This result indicates that when positive actions have a low probability. If the probability of sampling low-probability actions can be increased, the model will be optimized toward higher entropy.
Premise 2.
Positive samples from temperature-adjusted distributions induce predictable entropy change during training.
When positive samples are drawn from a higher-temperature version of the current policy distribution, training with these samples increases the model’s entropy, promoting exploration. Conversely, sampling from a lower-temperature distribution produces entropy reduction, guiding the model toward more deterministic behavior. Figure 2 illustrates this phenomenon: the entropy increases under high-temperature sampling and decreases under low-temperature sampling. Meanwhile, entropy increase is typically slower than entropy decrease, which aligns with intuition. This dynamic forms the empirical basis for AEPO’s temperature-controlled entropy feedback loop, enabling bidirectional regulation of entropy around a target value.
@@ 4.2 AEPO
Building upon the premises introduced above, AEPO is designed to achieve stable and controllable entropy regulation throughout RFT. Its core idea is to replace explicit entropy bonuses with a REINFORCE regularization term that adjusts the sampling distribution according to the current entropy state. In this way, AEPO modulates exploration implicitly—through data distribution—rather than directly through the loss function.
Formally, AEPO augments the standard GRPO objective with an additional policy gradient term applied to samples drawn from temperature-adjusted distributions. The overall policy gradient can be expressed as:
@@
Policy gradient as regularization. AEPO replaces conventional entropy regularization with a full policy gradient term that simultaneously enables entropy control and prevents entropy from dominating the optimization process. As illustrated in Fig. 3, entropy regularization alone tends to drive the optimization toward two extremes—either entropy collapse or entropy explosion. In the former case, the regularization term is too weak to reverse the monotonic entropy decay; in the latter, entropy becomes an irreversible dominant factor in optimization. By contrast, AEPO constrains entropy fluctuation within a narrow and stable range through the policy gradient mechanism, making it remarkably robust to hyperparameters.
@@
Distribution as regularization. AEPO regulates the optimization direction by adjusting the expected sampling distribution $\pi^{*}$ derived from the current policy. When the observed entropy $\mathcal{H}(\pi_{\theta_{\text{old}}})$ is below the target threshold $\mathcal{H}$, AEPO samples from the higher-temperature distribution $\pi^{T_{\text{high}}}_{\text{old}}$ to encourage exploration. Conversely, when $\mathcal{H}(\pi_{\theta_{\text{old}}})$ exceeds $\mathcal{H}$, AEPO samples from the lower-temperature distribution $\pi^{T_{\text{low}}}_{\text{old}}$ to promote stability. This bidirectional regulation mechanism achieves fine-grained entropy control, allowing the policy to maintain equilibrium between exploration and convergence.

### v6

# Arbitrary Entropy Policy Optimization Breaks The Exploration Bottleneck of Reinforcement Learning

@@ Abstract
Reinforcement Learning (RL) is essential for enhancing the reasoning capabilities of large language models (LLMs), yet the widely adopted Group Relative Policy Optimization (GRPO) suffers from entropy collapse, causing exploration to vanish and policies to converge prematurely. As a result, RL is widely believed to be incapable of expanding the reasoning frontier of LLMs. Existing entropy-regularized methods introduce an inevitable trade-off between reward and entropy, leading to exploration accompanied by non-negligible optimization bias. In this work, we prove that temperature-guided REINFORCE can modulate policy entropy, and propose Arbitrary Entropy Policy Optimization (AEPO), which reformulates entropy regularization as a policy-gradient optimization problem. Rather than manipulating entropy directly, AEPO implicitly regulates it by applying a REINFORCE regularization term on temperature-adjusted samples, ensuring that entropy is controlled but never dominates optimization, thereby enabling arbitrary and principled entropy regulation. Experiments show that AEPO outperforms RL baselines on both pass@1 and pass@$k$, and even surpasses the base model on pass@1024. By modulating entropy precisely, AEPO achieves more effective optimization dynamics and provides direct empirical evidence that entropy, exploration, and performance are intrinsically linked.
@@
Figure 1: Entropy across five runs of AEPO. By adjusting only the parameter $\mathcal{H}$, entropy can be controlled at different levels.
@@ 1 Introduction
Reinforcement Learning (RL) has become a cornerstone for enhancing the reasoning capabilities of large language models (LLM) (glm2024chat; touvron2023llama; schulman2017proximal; rafailov2023direct; zhong2024dpo; wang2024comprehensive). Entropy reflects the extent of exploration by measuring the model’s uncertainty and output diversity during reasoning (schulman2017equivalence; haarnoja2018soft; nachum2017bridging). Among existing approaches, Group Relative Policy Optimization (GRPO) has gained wide adoption due to its efficiency and scalability (shao2024deepseekmath; liu2024deepseek; guo2025deepseek). However, GRPO suffers from a well-documented drawback: entropy collapse: As training progresses, policy entropy declines monotonically, sampled outputs converge to nearly identical solutions, and the model prematurely adopts a deterministic policy with limited exploration (yu2025dapo; li2025disco; zhang2025edge). This limits the ability of RL to discover diverse reasoning strategies. Recent studies have shown that RL fails to broaden the reasoning capabilities of LLMs; instead, it merely sharpens behaviors within the base model’s existing knowledge without introducing genuinely new reasoning patterns (yue2025does). This highlights the necessity of effective exploration in RL for LLMs, and such exploration fundamentally requires breaking the entropy bottleneck (cui2025entropy).
@@
Although the problem of entropy collapse in GRPO has been repeatedly recognized, it remains fundamentally unsolved. Most existing variants incorporate entropy as an additional reward term in the gradient, such as entropy regularization (hou2025advancing; cui2025entropy; shen2025entropy) or advantage-weighted bonuses (cheng2025reasoning), to partially alleviate entropy collapse. However, these methods introduce an inevitable trade-off between reward and entropy, and thus bring non-negligible optimization bias when promoting exploration. In practice, entropy may oscillate between collapse and explosion instead of stabilizing in an optimal exploration regime, making the entropy–exploration–performance relationship hard to observe. As a result, it remains unclear whether entropy is a sufficient proxy for exploration and whether exploration itself consistently improves training outcomes. If significant performance improvement can be observed by adjusting entropy to a better range, it would indicate that exploration plays a crucial role in this process. If arbitrary entropy control can be achieved during RL, it would make it possible to realize exploration at any desired degree and, in turn, establish a principled connection among entropy, exploration, and performance.
Motivated by these challenges, we propose Arbitrary Entropy Policy Optimization (AEPO), which reformulates entropy regularization as a policy optimization, thereby fundamentally resolving entropy collapse. Instead of adding entropy bonuses, AEPO applies a REINFORCE policy gradient to temperature-adjusted samples, avoiding the unestimable bias introduced by conventional entropy-regularized methods. As shown in Fig. 1, AEPO maintains entropy oscillation around an arbitrary constant and is even theoretically capable of approximating arbitrary functions. AEPO achieves entropy control through three key design components:
  • Policy gradient as regularization: Instead of using an entropy bonus, AEPO applies a full policy-gradient term to samples with naturally high or low entropy, preventing entropy from dominating optimization while enabling stable exploration.
  • Temperature as regularization: Entropy is modulated through temperature-based sampling. When the current entropy falls below the target level, AEPO draws higher-temperature samples to increase entropy; when it exceeds the target, AEPO instead draws lower-temperature samples to reduce entropy.
  • REINFORCE as regularization: In RLVR, the REINFORCE algorithm can filter out negative samples without introducing bias, allowing positive samples to form a unidirectional gradient toward a better distribution.
@@
  • Breaking the entropy bottleneck: We propose Arbitrary Entropy Policy Optimization, which can stabilize entropy at arbitrary target levels, effectively eliminating entropy collapse in GRPO and enabling exploration beyond the longstanding entropy bottleneck. AEPO introduces a REINFORCE regularization mechanism that controls entropy without distorting the optimization objective, representing a fundamentally new paradigm for exploration.
  • Breaking exploration bottleneck: AEPO achieves consistent improvements over RL baselines on both pass@1 and pass@k, and even surpasses the base model on pass@1024, demonstrating that RL can conduct exploration rather than merely sharpening the base model’s knowledge.
@@ 2 Related Work
RL has become a central paradigm for post-training large language models (LLMs), aligning them with human feedback and task-specific objectives. Early methods, such as RLHF, leveraged policy optimization (e.g., PPO) to encode human preferences (openai2023gpt4; team2024gemini1_5; wei2023instructiongpt; liu2023visual), while Direct Preference Optimization (DPO) (rafailov2024direct) later improved efficiency by optimizing policies directly from preference data. Recent models like DeepSeek-R1 (guo2025deepseek), and Kimi-1.5 (team2025kimi1_5) extend RL through hybrid reward formulations and scalable optimization. Among them, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath; liu2024deepseek) has become the de facto baseline for reasoning-focused RL, yet suffers from entropy collapse that limits exploration of diverse reasoning strategies.
Entropy has long been regarded as a proxy for exploration in reinforcement optimization. Classical methods employ entropy regularization to stabilize training and encourage diversity (sutton1999policy; williams1992simple). More recent studies extend this idea to GRPO by introducing entropy bonuses into rewards or advantages (cheng2025reasoning; cui2025entropy; shen2025entropy). However, such approaches only yield coarse-grained effects: entropy still collapses as training proceeds, or the added bias distorts optimization.
@@ 3 Preliminary
Our work focuses on fine-tuning LLM using Reinforcement Learning with Verifiable Reward (RLVR), such as mathematical reasoning and code generation.
@@ 3.2 Entropy-regularization variants
Entropy-Adv(cheng2025reasoning) proposed an entropy-guided advantage shaping method. The key idea is to inject an entropy-based term into the advantage function during policy optimization. They define an entropy-based advantage term $\psi(\mathcal{H}_{t})$ and use it to shape advantage:
@@
$\nabla_{\theta}\mathcal{J}^{\text{shaped}}(\theta)=$ $\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta}(O|q)}$
@@ 4 Method
Arbitrary Entropy Policy Optimization (AEPO) is designed to achieve precise and stable control of policy entropy during RL. This section first introduces the key theoretical premises underlying AEPO’s design, which reveal the relationship between temperature, entropy, and policy updates, and then details the algorithmic formulation of AEPO.
$\begin{split}\mathcal{J}_{GRPO}(\theta)=&\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,\hat{A}_{i,t},\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,\hat{A}_{i,t}\Big],\\[5.69054pt] \mathcal{J}_{AEPO}(\theta)=&\mathcal{J}_{GRPO}+\alpha\;\mathbb{E}_{q\sim P(Q),\;\{o_{i}\}_{i=1}^{G}\sim\pi^{T}_{\theta_{\text{old}}}(O|q)}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\Big[r_{i,t}(\theta)\,R(q,o_{i}),\;\text{clip}\!\big(r_{i,t}(\theta),1-\epsilon,1+\epsilon\big)\,R(q,o_{i})\Big],\end{split}$ (1)
@@
4.1 Theoretical analysis
The design of AEPO is built upon two empirical premises that connect temperature-based sampling to entropy dynamics. These analyses establish the foundation for controllable entropy modulation without introducing bias into the optimization objective.
Lemma 4.1.
@@ 4.1 Theoretical analysis
Previous studies, such as du2025optimizing, show that increasing the sampling temperature generally broadens the model’s output distribution and raises its entropy. In the context of RL, temperature can adjust entropy during inference, but it does not modify the underlying policy itself. Simply increasing temperature to force diverse outputs leads to off-policy sampling, which undermines the consistency required for policy optimization. What is needed instead is a mechanism that directly adjusts the policy itself so that the original policy is capable of generating more diverse behaviors.
Assumption 4.2.
Assume that $T>1$ and the actor policy $\pi_{\theta}$ is a tabular softmax policy updated via the following equation:
$\theta_{k+1}=\theta_{k}+\eta\mathbb{E}_{a\sim\pi_{\theta_{k}}^{T}(\cdot|s)}[R(s,a)\cdot\nabla_{\theta}\text{log}\pi_{\theta_{k}}(a|s)].$
Theorem 4.3.
Denote the function
$\Delta H_{k}(T)=-\eta\cdot$ $\sum_{a^{*}\in\mathcal{A}^{*}}\pi_{\theta_{k}}^{T}(a^{*}|s)\times$
$\text{Cov}_{\pi_{\theta_{k}}}(z_{k}(s,a^{\prime}),\delta_{a^{\prime},a^{*}}-\pi_{\theta_{k}}(a^{\prime}|s)).$
Then we have
$\Delta H_{k}(T)\approx H(\theta_{k+1})-H(\theta_{k}),$
and
$\Delta H_{k}^{{}^{\prime}}(T)|_{T=1}>0\quad\text{if}\quad\Delta H_{k}(T)|_{T=1}<0.$
The proof is provided in Appendix B.5. In more intuitive terms, the preceding theorem implies that if the policy entropy begins to collapse during training, we can counteract this downward trend by performing updates using REINFORCE policy gradient of high-temperature samples.
Corollary 4.4.
High-temperature REINFORCE induces a relative increase in policy entropy, while low-temperature REINFORCE induces a relative decrease in policy entropy.
Within the RLVR framework, the binary nature of the reward signal gives rise to an inherent filtering mechanism for negative samples. Building upon this property and the preceding premises, we can formulate a clear insight: computing the policy gradient with temperature-adjusted samples enables predictable and controllable entropy modulation. As illustrated in Figure 2, entropy increases under high-temperature sampling and decreases under low-temperature sampling, with the increase typically occurring more gradually than the decrease. This dynamic forms the empirical foundation for AEPO’s temperature-controlled entropy feedback loop, allowing bidirectional regulation of entropy around a target value.
Figure 2: Entropy dynamics under temperature-controlled REINFORCE. High-temperature REINFORCE increases entropy, promoting exploration, while low-temperature REINFORCE reduces entropy.
@@ 4.2 AEPO
Building on the theoretical premises described above, AEPO achieves stable and controllable entropy regulation by replacing explicit entropy bonuses with a REINFORCE-based regularization mechanism that adjusts the sampling temperature according to the current entropy state. Concretely, AEPO augments the GRPO objective (without KL divergence) with an additional policy gradient term applied to temperature-adjusted samples. The resulting policy gradient is given by:
@@
Policy gradient as regularization. AEPO replaces conventional entropy regularization with a full policy gradient term that simultaneously enables entropy control and prevents entropy from dominating the optimization process. As illustrated in Fig. 3, entropy regularization tends to drive the optimization toward two extremes—either entropy collapse or entropy explosion. In the former case, the regularization term is too weak to reverse the monotonic entropy decay; in the latter, entropy becomes an irreversible dominant factor in optimization. By contrast, AEPO constrains entropy fluctuation within a narrow and stable range through the policy gradient mechanism, making it remarkably robust to hyperparameters.
@@
Temperature as regularization. AEPO regulates the optimization direction by adjusting the sampling temperature. When the observed entropy $\mathcal{H}(\pi_{\theta_{\text{old}}})$ is below the target threshold $\mathcal{H}$, AEPO samples from the higher-temperature distribution $\pi^{T_{\text{high}}}_{\text{old}}$ to encourage exploration. Conversely, when $\mathcal{H}(\pi_{\theta_{\text{old}}})$ exceeds $\mathcal{H}$, AEPO samples from the lower-temperature distribution $\pi^{T_{\text{low}}}_{\text{old}}$ to promote stability. This bidirectional regulation mechanism achieves fine-grained entropy control, allowing the policy to maintain equilibrium between exploration and convergence.